Prove that x(1 + x) > (1 + x) In(1 + x) > x

1
Expert's answer
2022-03-24T03:12:03-0400

"x<(1+x)ln(1+x)<x(1+x)"

"let f(x)=(1+x)ln(1+x)-x..............(1)"

"f'(x)=(1+x)(\frac{1}{1+x})+ln(1+x)1-1"


"f'(x)=1+ln(1+x)-1=ln(1+x)>0"


i.e. "x>0; f(x)>f(0), where f(0)=(1+0)ln(1+x)-0=0"

Hence, f(x)>0

We use 1 above

"(1+x)ln(1+x)-x>0 ;"


"(1+x)ln(1+x)>x.....(M)"

Let g(x)=x(1+x)-(1+x)ln(1+x)........(2)

"g'(x)=(2x+1)-((1+x)\frac{1}{1+x}+ln(1+x)1"


"=2x+ln(1+x)>0"

From this we find that,

"g(x)>g(0), where \,g(0)=0(1+0)-(1+0)ln(1+0)=0"

g(x)>0...use (2)

"x(1+x)-(1+x)ln(1+x)>0"

"x(1+x)>(1+x)ln(1+x).........(N)"


From (N)and (M)

x<(1+x)ln(1+x)<x(1+x)

Hence it's clear that,


"x(1+x)>(1+x)ln(1+x)>x"